OpenStudy (anonymous):

Alex’s gasoline consumption, g, varies inversely with the price of a gallon, p. Alex uses 22 gallons of gas per week when the price is $3.49 per gallon. How many gallons will he use if the price goes up to $3.89 per gallon? Round to the nearest gallon.

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hi, my name is Stephen. \[\frac{22g * 3.49}{gallon} = \frac{x gallons * 3.89}{gallon} \] (as the price goes up, his usage goes down...) X = 22*3.49/3.89 = 19.74 gallons \[x = \frac{22 * 3.49}{3.89} = 19.74 gallons\] Now, it's up to you to round it to the tenth.